3|x-2|=18
Solve for x

Answer by xdragonfight(116)

THERE IS 2 SOLUTION !!
3x-6=18 and 3x-6=-18
3x-6=18
+6 +6
3x=24
24/3=8
x=8
ok here 2nd solution
3x-6=-18
+6 +6
3x=-12
x=-4
so u will put ur answer as -4 and 8